OpenStudy (decentnabeel):

Suppose a book has 100 names listed book has 200 pages,how many books are 0r 200 million?

OpenStudy (decentnabeel):

@mww @AloneS @ganeshie8

OpenStudy (decentnabeel):

Suppose a book has 100 names listed each book has 200 pages,how many books are 0r 200 million?

OpenStudy (decentnabeel):

@pooja195 @satellite73

OpenStudy (decentnabeel):

@KendrickLamar2014

OpenStudy (legomyego180):

this isnt a complete question

OpenStudy (mrnood):

is this supposed to be: 100 names per page 200 pages per book How many books for 200 million names? If so It would be easy to do this using exponent notation each book contains 1*10^2 names per page there are 2 * 10^2 pages multiply those to get 'names per book' you want 'total names' =200 *10^6 (= 2 *10^8) so divide the 'total names'/ 'names per book' this is very easy using exponents.